OpenStudy (anonymous):

Choose the correct simplification of 5^4\5^2

OpenStudy (unklerhaukus):

\[\frac{5^4}{5^2}=\frac{5\times5\times5\times5}{5\times5}=5^{(4-2)}\]

OpenStudy (anonymous):

Possible Answers: a. 10 b. 1/25 c. 1/10 d. 25

OpenStudy (anonymous):

(5^2)^2/5^2 5^2=25